Plain M&Ms candies come in six colors: yellow, red, orange, brown, blue, and green. The makers of M&Ms say that 20% of the
Answer #1

Proportion of brown candies = 0.15

P(yellow or brown) = 0.20+0.15 = 0.35

Proportion of green = 1 - (0.20+0.20+0.20+0.15+0.15) = 0.10

P(red or orange) = 0.20+0.20 = 0.40

P(not red or orange) = 1 - 0.40 = 0.60

Correct statement is:

If we randomly select one candy at random from all plain M&M's that are produced, there is a 60% chance that candy will not be red or orange.